### v2

# Learning Transferable Friction Models and LuGre Identification via Physics-Informed Neural Networks

Abstract
Accurately modeling friction in robotics remains a core challenge, as robotics simulators like MuJoCo and PyBullet use simplified friction models or heuristics to balance computational efficiency with accuracy, where these simplifications and approximations can lead to substantial differences between simulated and physical performance. In this paper, we present a physics-informed friction estimation framework that enables the integration of well-established friction models with learnable components, requiring only minimal, generic measurement data. Our approach enforces physical consistency yet retains the flexibility to capture complex friction phenomena. We demonstrate, on an underactuated and nonlinear system, that the learned friction models, trained solely on small and noisy datasets, accurately reproduce dynamic friction properties with significantly higher fidelity than the simplified models commonly used in robotics simulators. Crucially, we show that our approach enables the learned models to be transferable to systems they are not trained on. This ability to generalize across multiple systems streamlines friction modeling for complex, underactuated tasks, offering a scalable and interpretable path toward improving friction model accuracy in robotics and control.
I Introduction
Machine learning and deep learning have gained momentum in solving complex problems in computer vision, natural language processing, and generative modeling. Large, high-quality datasets are often available in these domains or can be synthesized to train data-hungry models. For instance, virtual 3D environments provide large synthetic data for tasks like obstacle detection, mapping, and navigation in robotics [21, 22]. These data-driven strategies work well in purely digital domains or when high-fidelity virtual environments are available.
However, in real-world physical interactions involving contact and friction, high-quality data is scarce. Real-world experiments are expensive, time-consuming, and subject to noise or wear-and-tear constraints that make large-scale data collection impractical. This challenge is amplified when robotic systems must deal with unpredictable or varying contact conditions, such as friction changes or impacts during locomotion and manipulation. In particular, underactuated systems that rely on friction for locomotion face challenges in modeling continuous sliding contacts accurately [7]. Many robotics simulators, such as MuJoCo or PyBullet, compensate by using simplified friction models or heuristics to balance computational efficiency with accuracy [12]. Since friction is highly dependent on local surface properties, velocity, and normal force, these simplifications can lead to substantial differences between simulated and physical performance. This mismatch, commonly referred to as the sim-to-real gap, can be especially pronounced in applications where friction plays a key role in system stability and control. To mitigate the resulting inaccuracies, methods such as Model Predictive Control (MPC) incorporate high-frequency feedback by repeatedly solving Optimal Control problems at rates up to 1 kHz [6, 15, 10, 5].
I-A Related Work
Physics-Informed Neural Networks (PINNs) have found applications across a wide range of fields, including fluid dynamics, plasma physics, quantum chemistry, and material science [2, 14, 17, 23, 9]. PINNs have shown particular promise in friction modeling, structural hysteresis prediction, and robotic joint dynamics [11, 4, 24]. They offer significant advantages such as data efficiency, interpretability, and computational efficiency, reducing dependency on extensive datasets and discretizations [18, 8, 11]. However, training complexity, scalability, and generalization remain open challenges, motivating ongoing research on hybrid models and optimization strategies [4, 8, 11].
A few researchers have proposed learning-based friction models to bridge the sim-to-real gap. Sorrentino et al. incorporate friction data a priori into the training process, using measured friction forces to guide training losses [24]. Others omit direct friction force measurements but rely on simplified linear friction terms [20, 16], which may be too restrictive to capture complex frictional phenomena, particularly for surface friction characteristics such as stick-slip behavior under time-varying normal forces or in underactuated settings. Moreover, large-scale data collection for frictional interactions is far more demanding than collecting images or text, limiting how effectively purely data-driven methods can scale.
More broadly, hybrid approaches that combine physics-based models with data-driven components have gained traction across dynamical systems modeling [8]. These methods range from using neural networks to learn residual corrections on top of known dynamics to embedding physical structure directly into the learning architecture. Our work falls into the latter category: rather than learning a correction term or using physics only as a regularizer, we embed the LuGre friction model structure into the network’s forward pass, enabling the framework to learn both the friction behavior and the underlying model parameters from state data alone.
Decades of friction modeling research have yielded detailed models and theoretical insights, most notably advanced friction representations such as the LuGre model [3], which captures dynamic friction behaviors like stick-slip motion and the Stribeck effect. However, a hybrid approach that integrates well-established friction models into a learning-based framework remains underexplored. Our hybrid approach enables the learned friction models to gain interpretability and physical consistency while still leveraging data to correct for model inaccuracies in the face of real-world complexities.
I-B Aim of the Paper
We propose a physics-informed learning approach that learns a friction model with minimal reliance on extensive friction datasets and incorporates learnings from prior research on friction modeling. Our methods use only the system’s states and the governing equations of motion (EoMs). Building on ideas from PINNs, we embed the friction model within a loss term based on equations of motion, ensuring that friction estimates remain consistent with the physics that governs the system. This approach is versatile since it can be implemented as a black-box neural network or as a parameter estimation model that explicitly identifies the parameters of a known friction formulation. Rather than starting from scratch, our hybrid formulation leverages decades of research in friction models and provides additional benefits such as interpretability and potential parameter/model reuse across different systems with similar surface contact properties.
Our framework differs from standard PINN formulations in several key aspects. We drop the data loss term entirely: the network is trained solely through the EoM residual, with no direct supervision on the output quantity. The network learns to predict quantities that are never directly measured in the training data, namely friction force in the black-box models and the LuGre internal bristle state $z$ in the parameter estimation models. The physics loss is therefore not a regularizer augmenting data-driven learning, but the exclusive training signal for inferring unobserved physical quantities from readily available state measurements. Furthermore, in our parameter estimation models, the LuGre friction equations are architecturally embedded in the network’s forward pass: the LuGre parameters ($\sigma_{0,1,2}$, $\mu_{c,s}$, $v_{s}$) are learned as trainable variables, and the friction force is computed through the LuGre structure rather than approximated by a generic function. This structural integration is reinforced by a consistency loss $\mathcal{L}_{\dot{z}}$ that couples the network’s output, obtained via automatic differentiation, to the LuGre state evolution equation, ensuring that the learned internal state and parameters remain mutually consistent.
In contrast to prior learning-based friction methods that either require direct friction force measurements during training [24] or rely on simplified linear friction terms [20, 16], our approach requires neither, learning complex dynamic friction behavior from generic state data alone. To the best of our knowledge, this is one of the few frameworks providing PINN-based transferable learned friction models that focus on surface friction rather than joint or drive friction. We demonstrate that models trained on one system generalize to a different dynamical system sharing the same contact surface, and that the parameter estimation variant identifies LuGre parameters with accuracy comparable to established methods such as Nelder–Mead and genetic algorithms while being significantly faster. Our aim is to show that even with limited and noisy datasets, high-fidelity friction models can be efficiently learned, offering improved friction model accuracy over the simplified models commonly used in robotics simulators.
The rest of the paper is structured as follows. In Section II, we outline the preliminaries in the form of the LuGre friction model and PINNs that are necessary for our framework. Section III outlines our PINN-based framework. Section IV shows that our framework is suitable for learning friction models that can be used as in-simulation friction models and online friction estimators, as well as being transferable to different dynamical systems.
II Preliminaries
II-A LuGre Friction Model
Introduced by Canudas de Wit et al. (1995), the LuGre friction model significantly advanced friction modeling by addressing the shortcomings of classical static friction models like Coulomb and viscous friction, particularly at low velocities and during velocity reversals [3].
Traditional models often neglect dynamic behaviors such as hysteresis and the Stribeck effect, where friction force decreases after a certain velocity threshold [13]. The LuGre model captures these effects by introducing an internal state representing microscopic bristle deflection at the frictional interface. This state evolves with relative velocity, allowing simulation of pre-sliding displacement and varying break-away forces [1].
TABLE I: Parameters of LuGre friction model
Parameter Description
$\boldsymbol{\sigma_{0}}$ Bristle stiffness
$\boldsymbol{\sigma_{1}}$ Bristle damping
$\boldsymbol{\sigma_{2}}$ Viscous damping coefficient
$\boldsymbol{F_{s}},\boldsymbol{\mu_{s}}$ Static friction force and coefficient
$\boldsymbol{F_{c}},\boldsymbol{\mu_{c}}$ Coulomb friction force and coefficient
$\boldsymbol{v_{s}}$ Stribeck velocity
$\boldsymbol{\alpha}$ Transition shape factor
The LuGre friction model is characterized by the internal state evolution (1) and the linear friction force (2) shown below:
$\dot{z}$ $=v-\frac{\sigma_{0}|v|}{F_{c}+(F_{s}-F_{c})e^{-(|v|/v_{s})^{\alpha}}}z$ (1)
$F_{f}$ $=\sigma_{0}z+\sigma_{1}\dot{z}+\sigma_{2}v$ (2)
Here, $z$ represents the internal state variable corresponding to the average deflection of microscopic bristles at the contact interface, and $v$ denotes the relative velocity between the surfaces. The denominator of the function that is multiplied by $z$ in (1) encapsulates the velocity-dependent nonlinear characteristics of friction, notably the Stribeck effect, which describes the reduction in friction force with increasing velocity past a certain threshold. In this paper, we represent Coulomb and static friction forces $F_{c}$ and $F_{s}$ in (1) as functions of their respective friction coefficients $\mu_{c}$, $\mu_{s}$ and the normal force $F_{N}$. $F_{f}$ in (2) is the resulting friction force, where the LuGre friction parameters are defined in Table I.
$F_{c}$ $=\mu_{c}|F_{N}|$ (3)
$F_{s}$ $=\mu_{s}|F_{N}|$ (4)
These parameters are typically determined through experimental identification methods, which involve measuring frictional forces under controlled conditions to fit the model accurately to observed behavior [25]. By appropriately selecting and calibrating these parameters, the LuGre model can effectively simulate various frictional phenomena, including stick-slip motion, hysteresis, and pre-sliding displacement, making it a valuable tool in the analysis and control of mechanical systems subject to friction [1].
Figure 1: Pendulum-on-a-Box system
The LuGre friction model represents a foundational tool in the study and control of frictional systems, offering a dynamic framework that captures a wide range of frictional behaviors beyond the capabilities of traditional static models. Although some challenges, including drift and discrepancies in non-stationary regimes, continue to drive refinements and new friction models [19], for the purpose of this paper, we will be focusing on the base LuGre model defined by (1) and (2).
II-B Physics-Informed Neural Networks (PINNs)
Physics-Informed Neural Networks (PINNs) introduced by Raissi et al. (2019) integrate physical laws described by partial differential equations (PDEs) directly into neural network frameworks and were proposed as a framework for solving forward and inverse problems involving nonlinear PDEs [18]. Unlike traditional machine learning approaches, PINNs leverage the governing equations of physical phenomena as constraints during the training process, enhancing model interpretability, accuracy, and efficiency.
A typical PINN formulation involves embedding PDEs, initial conditions, and boundary conditions into the loss function of the neural network. The loss function comprises two components: Data Loss and Physics Loss, where the former ensures the network output aligns with available observational data utilizing a mean square error (MSE) between the networks’ predictions and observational data, and the latter penalizes deviations from the governing physical equations. PINNs thus aim to minimize a composite loss function:
$\mathcal{L}=\mathcal{L}_{\mathcal{D}}+\mathcal{L}_{\mathcal{P}}$ (5)
where $\mathcal{L}_{\mathcal{D}}$ and $\mathcal{L}_{\mathcal{P}}$ denote the data loss and the physics loss respectively.
TABLE II: Parameters of the PoB System
Parameter Description Value/Unit
$m_{b}$ Mass of the box 0.5 kg
$m_{L}$ Mass of the link 1 kg
$L$ Length of the link 0.5 m
$d$ Distance from pivot to link CoM 0.25 m
[…]
II-C Pendulum-on-a-Box System
We introduce the pendulum-on-a-box (PoB) system in Fig. 1, which is a modification of the cart-pole system where the pole is powered instead of the cart, and the cart wheels are removed. The system must leverage surface friction by swinging the arm to achieve locomotion. We chose this system as a testbed for our framework since it shows properties similar to more challenging problems in robotics and locomotion. The system is non-linear and underactuated, and the normal force constantly fluctuates during phases of locomotion, while friction is paramount in achieving locomotion. The Lagrangian method is used to derive the equations of motion (6)-(8) and put into manipulator equations of the form in (9).
Figure 2: I/O of Blackbox Models (a), Parameter Estimation Models (b) and corresponding loss terms. Automatic differentiation (AD) is used to compute $\hat{\dot{z}}_{model}$
$(m_{b}+m_{L})\ddot{x}_{b}=-m_{L}d(\ddot{\theta}\cos\theta-\dot{\theta}^{2}\sin\theta)-F_{f}$ (6)
$(m_{b}+m_{L})(\ddot{y}_{b}+g)=m_{L}d(\ddot{\theta}\sin\theta+\dot{\theta}^{2}\cos\theta)\hfill$ (7)
$(J_{L}+m_{L}d)\ddot{\theta}=-m_{L}d\big((\ddot{x}_{b}-\dot{y}_{b}\dot{\theta}+\dot{y}_{b})\cos\theta$-
$\hskip 95.0pt(\dot{x}_{b}\dot{\theta}+\ddot{y}_{b}+\dot{x}_{b}-g)\sin\theta\big)+\tau$ (8)
where the states $x_{b}$, $y_{b}$, $\theta$ are defined as the x and y position of the box and the angle of the link in radians, respectively.
$\textbf{M}(\textbf{q})\ddot{\textbf{q}}+\textbf{C}(\textbf{q},\dot{\textbf{q}})\dot{\textbf{q}}+\textbf{T}_{g}(\textbf{q})g=\textbf{B}\textbf{u}$ (9)
where, $\textbf{q}=[x_{b},y_{b},\theta,\dot{x}_{b},\dot{y}_{b},\dot{\theta}]^{T}$, $\textbf{u}=[-F_{f},0,\tau]^{T}$, updating our definition of $\boldsymbol{\dot{z}}$ to (10).
$\dot{z}=\dot{x}_{b}-\frac{\sigma_{0}|\dot{x}_{b}|}{F_{c}+(F_{s}-F_{c})e^{-(|\dot{x}_{b}|/v_{s})^{\alpha}}}z$ (10)
The updated $\dot{z}$} is then used to calculate the friction force using (2) and is incorporated into the equations of motion to simulate the system. The numerical values of the parameters for the PoB system are listed in Table II. Data generation for training using this system will be described in more detail in the next section.
III Methods
We propose two different neural networks for learning generalizable friction models without a priori knowledge of friction characteristics in the data. Henceforth, for our PINN approach, the data loss term $\mathcal{L}_{\mathcal{D}}$ in (5) is dropped. Our first approach is a blackbox (BB) friction model in Fig. 2a that estimates the friction force directly, and the second is a parameter estimation (PE) model in Fig. 2b that estimates the internal LuGre state, $z$, along with the LuGre parameters. We use feed-forward neural networks with fully connected layers for all approaches.
III-A PINNs for Learning Blackbox Friction Estimators
Our approach is based on PINNs to learn generalizable friction models. The friction network BB1 in Fig. 2(a, top), as inputs, takes the relative velocity between surfaces $\dot{x}_{b}$} and the absolute value of the normal force estimate $\hat{F}_{N}$}. Given that directly measuring the normal force is not always practical, we utilize the equations of motion to estimate the normal force based on the system’s states as outlined in (11) with the assumption that the system does not break contact with the surface.
$\hat{F}_{N}=m_{L}d\dot{\theta}^{2}\cos\theta-(m_{b}+m_{L})g$ (11)
The physics loss function is based on (6) in the equations of motion for the PoB system.
[…]
This initial setup is sufficient to be used as friction models in simulation for any system. However, we also introduce an additional neural network BB2 in Fig. 2(a, bottom) to enable our approach to be feasible for online friction estimation. We provide an additional input to the neural network that is defined as the “would be acceleration” of the contact point if no friction was present. In the PoB system, this equates to the would-be box acceleration in zero friction conditions, which is defined by (13).
$\ddot{x}_{b}^{*}=m_{L}d(-\ddot{\theta}\cos(\theta)+\dot{\theta}^{2}\sin\theta)/(m_{b}+m_{L})$ (13)
This input is derived from the system states and known EoMs and provides the network with the necessary information for online friction estimation. Inputs and outputs of these networks are outlined in Fig. 2a.
III-B PINNs for LuGre State and Parameter Estimation
Our second approach learns the internal state of the LuGre friction model and the underlying parameters that make it up. The physics loss in this approach includes the estimated friction force in the form of LuGre friction and an additional term $\mathcal{L}_{\dot{z}}$ with a scaling factor $\lambda$ that is used to drive the learning towards $\hat{z}$ and $\hat{\dot{z}}_{\text{model}}$ terms that are consistent with the LuGre structure.
$\mathcal{L}$ $=\mathcal{L}_{\mathcal{P}}+\lambda\mathcal{L}_{\dot{z}}$ (14)
[…]
$\hat{F}_{f}$ $=\hat{\sigma}_{0}\hat{z}+\hat{\sigma}_{1}\hat{\dot{z}}_{\text{model}}+\hat{\sigma}_{2}\dot{x}_{b}$ (16)
$\hat{F}_{c}$ $=\hat{\mu}_{c}|\hat{F}_{N}|$ (17)
$\hat{F}_{s}$ $=\hat{\mu}_{s}|\hat{F}_{N}|$ (18)
$\hat{\dot{z}}_{\text{LuGre}}$ $=\dot{x}_{b}-\frac{\hat{\sigma}_{0}|\dot{x}_{b}|}{\hat{F}_{c}+(\hat{F}_{s}-\hat{F}_{c})e^{-(|\dot{x}_{b}|/\hat{v}_{s})^{2}}}\hat{z}$ (19)
In this approach, the neural network PE1 in Fig. 2(b, top) has output $\hat{z}$ and $\hat{\dot{z}}_{\text{model}}$ denotes its derivative. $\hat{\sigma}_{0,1,2}$, $\hat{\mu}_{c,s}$ and $\hat{v}_{s}$ are estimated parameters of the LuGre friction model; $\alpha$} in (1) is set to 2 as it is commonly adopted. $\hat{\dot{z}}_{\text{LuGre}}$ is an estimate of $\dot{z}$ from LuGre formulation in (10) using the $\hat{z}$, and estimated LuGre parameters. The LuGre parameters to be estimated are added to the neural network as trainable variables to be learned and adjusted during training. Similar to the blackbox networks, $\ddot{x}_{b}^{*}$ is used as an additional input for PE2 in the secondary approach in Fig. 2(b, bottom) for parameter estimation models to enable online friction estimation where necessary.
Figure 3: BB1 and PE1 friction estimations for: (a) Traj. 1, (b) Traj. 2. Overall MSE (N2) for in-simulation / online: (a) BB1: 0.041 / 0.124, PE1: 0.014 / 0.032; (b) BB1: 0.092 / 1.674, PE1: 0.012 / 1.591.
Figure 4: Online estimation performances of BB1,2 and PE1,2 for Traj. 2. Overall MSE (N2): BB1: 1.674, BB2: 0.174, PE1: 1.591, PE2: 0.066.
III-C Data Generation and Training
Training data for the PINN is generated in MATLAB R2023a by simulating the system dynamics with ode45. The LuGre model in (2) and (10) is used as ground truth due to its complexity and broad adoption.
Six trials were collected at 400 Hz, yielding about 5800 samples over 14.5 s. Five trials are pendulum swings at the same excitation frequency with amplitudes ranging from $\pm 35^{\circ}$ to $\pm 65^{\circ}$, each lasting 2 s. The sixth trial is a PoB translation in +x+x using a custom swing trajectory. Gaussian noise at 5% of each signal’s standard deviation is added to simulate sensor noise. These short, noisy trajectories are chosen to demonstrate learning from minimal generic data. Networks are trained on the noisy set using TensorFlow 2.15.0 in Python 3.10.12. Architectures are as follows: BB1 and PE1 use four hidden layers with 128 neurons per layer, and BB2 and PE2 use four hidden layers with 512 neurons per layer.
All networks use the Adam optimizer with ReLU activations, and are trained on the full batch for 10,000 epochs. BB1 and PE1 use a fixed learning rate of $1\!\times\!10^{-4}$. BB2 and PE2 use an initial learning rate of $1\!\times\!10^{-3}$ with adaptive learning rate scheduling that halves the rate when the loss plateaus, down to a minimum of $1\!\times\!10^{-5}$. For the parameter estimation models PE1,2, the scaling factor in (14) is set to $\lambda=1\!\times\!10^{5}$ to balance the EoM residual $\mathcal{L}_{\mathcal{P}}$ against the bristle deflection rate consistency loss $\mathcal{L}_{\dot{z}}$, which operate at different magnitudes.
Figure 5: Friction force characteristics for (a) PE1, (b) BB1, (c) LuGre model
Figure 6: Spring-Damper on a Box (SDoB) system
IV Results and Discussion
We evaluate the proposed friction estimators for in-simulation use, online estimation, transferability across different dynamical systems, and LuGre parameter identification speed and accuracy.
TABLE III: Estimated LuGre Friction Parameters
Pars. Ground Truth PE1 PE2 Nelder–Mead Genetic Alg. Nonlinear LS
$\sigma_{0}$ (N/m) $1.00\!\times\!10^{5}$ $1.20\!\times\!10^{5}$ $1.02\!\times\!10^{5}$ $1.03\!\times\!10^{5}$ $1.04\!\times\!10^{5}$ $99.9\!\times\!10^{5}$
$\sigma_{1}$ (Ns/m) 316.23 346.41 319.37 320.36 322.43 316.21
$\sigma_{2}$ (Ns/m) 0.40 0.41 0.44 0.20 0.47 1
$\mu_{c}$ 0.30 0.28 0.30 0.30 0.30 0.23
$\mu_{s}$ 0.60 0.60 0.61 0.49 0.59 0.47
$v_{s}$ (m/s) $10.0\!\times\!10^{-4}$ $8.00\!\times\!10^{-4}$ $9.96\!\times\!10^{-4}$ $4.13\!\times\!10^{-4}$ $11.9\!\times\!10^{-4}$ $57.9\!\times\!10^{-4}$
$t_{comp}$ (min) - 9.46 10.54 15.18 67.80 24.14
IV-A PINN Friction Models for Use in Simulation
We evaluate BB1 and PE1 models as in-simulation (dashed) friction models and online friction estimators (solid) in Fig. 3. In the first approach, the friction force estimates from the trained neural networks are incorporated into (6), and the PoB system is simulated using the EoMs outlined in (6)-(8) and the differential equations are solved using ode45 in MATLAB. In the second approach, the data collection is done with the ground truth LuGre friction model, and the noisy data at each timestep is fed into the trained neural networks to estimate the friction force at that time step for online estimation of the friction force. We test the models on two different trajectories: a constant oscillation of the pendulum at a set range and a custom trajectory designed to move the PoB system in the +x direction by swinging the pendulum at different speeds to leverage the stick and slip nature of friction between the surface and the PoB system. For the sake of simplicity, in the rest of this paper we will refer to these trajectories as “Traj. 1” and “Traj. 2” respectively.
The trained friction models BB1 and PE1 appear to perform well in both in-simulation and online estimation for Traj. 1 in Fig. 3a. However, in trajectories similar to Traj. 2 (Fig. 3b), where there are extended stationary periods during which the friction force depends on the internal forces of the system, their online estimation capabilities fall short. In these trajectories, the inputs to the models BB1 and PE1 ($\dot{x}_{b}$$F_{N}$) do not provide sufficient information for estimating friction and therefore fail in the stationary regime when used as online estimators. PE1 achieves the lowest in-simulation MSE across both trajectories (0.014 and 0.012 N2), while online estimation errors on Traj. 2 are significantly higher (BB1: 1.674, PE1: 1.591 N2) due to insufficient information during stationary periods.
Dynamic friction models like LuGre rely on micro-motions after each numerical integration step to deflect the bristles and generate a reactive force, even when the object appears stationary. Our trained models BB1 and PE1 similarly capture these phenomena when used in simulation, but this mechanism is unavailable during online estimation. We see this as a limitation and propose a secondary approach for online estimation using PINNs.
IV-B PINN Friction Models for Online Friction Estimation
Friction estimation is an important part of planning trajectories and tracking accuracy; hence, our proposed secondary models, BB2 and PE2, are improvements upon BB1 and PE1 in that they can be used to estimate friction on the go. The additional input of the secondary models enables them to break free of the limitations that BB1 and PE1 suffer from by having only velocity and normal force information and removes the need to simulate the dynamics for reliable friction estimation. Fig. 4 compares the online estimation performance of all the PINN-based friction estimation models. The 3-input models substantially reduce online estimation error on Traj. 2: BB2 and PE2 achieve MSEs of 0.174 and 0.066 N2 (reductions of 90% and 96% over their 2-input counterparts).
Figure 7: Models trained on PoB for friction estimation in the SDoB system.
IV-C Transferability of Learned Friction Models and Limitations
Our approach enables the learned models to be transferable to different systems that are deployed in the same environment. Learned models exhibit behavior similar to the LuGre model for different velocity and normal force pairs (Fig. 5), enabling them to be used on different systems than the ones they are trained on if the system is deployed in the same environment. To test this, we introduce a system called Spring-Damper on a Box (SDoB) illustrated in Fig. 6, consisting of two masses connected by a spring and a damper pulled by an external force applied on the bottom mass ($m_{1}$). The top mass ($m_{2}$) in this system is free to move in the y direction, which results in varying normal force throughout the overall system trajectory. We simulated the SDoB system using the same underlying LuGre friction model as the PoB system and tested the learned friction models that are trained on the PoB system to estimate friction on the SDoB system (Fig. 7). Our results strongly suggest that the training framework proposed in this paper enables the trained models to be transferable to different systems or to the same dynamic system with different parameters. However, the extent of this transferability is inherently limited: significant modifications to the underlying system properties, such as changes in stiffness or flexibility, or the introduction of new materials at the contact interface with different frictional characteristics, may reduce model accuracy and require re-training or adaptation.
Figure 8: Comparison of various parameter estimation methods with respect to computation time and mean squared error (MSE) of friction force using the estimated parameters.
IV-D Parameter Estimation Performance
Choosing an appropriate friction model for each setup is an important decision. The complexity of the friction model, the identification process for parameters, the time it takes, and the computational cost are all essential decision factors in deciding which friction model is the best for a particular setup. Although the LuGre friction model accurately captures most aspects of the actual frictional behavior, such as stick-slip, it requires more work to identify the parameters that fit the experimental data.
We show that our methods employing PINNs enable faster identification of friction models of comparable complexity with sufficient accuracy. Table III compares the parameter estimation performance of established methods with our two models $PE_{1,2}$ shown in Fig. 2b. The performance of our models is on par with the other methods. A performance trade-off between mean squared error (MSE) of the estimated friction forces and computation time is illustrated in Fig. 8. While our approach may not replace full parameter identification in safety-critical systems where maximum accuracy is essential, it provides a faster alternative for estimating LuGre parameters. Although Nelder-Mead simplex method appears competitive in Fig. 8, its parameter estimation accuracy was worse (Table III).
V Conclusion and Future Work
We formulated a PINN-based friction estimation framework for creating transferable learned friction models. We demonstrated the efficacy of these models both as in-simulation friction models and online friction estimators, showed their transferability to a dynamical system they were not trained on, and compared the framework’s computational speed and accuracy in LuGre parameter identification against established methods.
The present results are validated using synthetic data generated from the LuGre model. While this demonstrates that the framework can recover complex friction behavior from minimal state measurements, experimental validation on physical hardware is needed to fully assess its applicability to real-world systems. Further work includes extending the framework to multi-contact and higher degree-of-freedom systems, systematic comparison with other hybrid learning-based friction modeling approaches, and integration with existing control methods and trajectory optimization. Expanding the framework to accommodate online adaptation to changing friction properties is also planned. In each such scenario, we anticipate a Pareto trade-off of the same general form (Fig. 8), with the exact trade-offs between computational speed and accuracy of motion being a key concern.